OpenStudy (anonymous):

Solve negative two-thirds times a plus two-fifths equals three-fourths for a.

OpenStudy (imstuck):

\[-\frac{ 2 }{ 3 }a+\frac{ 2 }{ 5 }=\frac{ 3 }{ 4 }\]

OpenStudy (imstuck):

Is that it?

OpenStudy (anonymous):

yes

OpenStudy (imstuck):

Ok you need to eliminate all the denominators by finding the LCM of all the denominators you have. The LCM between 3 and 4 and 5 is 60. There is no lower number that all 3 go into. So you multiply each term by 60, like this:

OpenStudy (imstuck):

\[(60)-\frac{ 2 }{ 3 }+(60)\frac{ 2 }{ 5 }=(60)\frac{ 3 }{ 4 }\]

OpenStudy (imstuck):

It will be easier to do it like this.